# Wall crossing and the Fourier–Mukai transform for Grassmann flops

### v1

@@ 4.2. Products of projective bundles
In this case for any $j$ we have the relationship between Gamma functions:
@@
$\hat{\Gamma}(TX_{T}/B)=\prod_{k=1}^{r}\prod_{j=1}^{n}\left(\hat{\Gamma}(1+(-y_{k}-x_{j}))\Gamma(1+(y_{k}+z_{j}))\right).$
@@
$H^{B}_{T}$ $:=(2\pi\mathbbm{i})^{-\deg_{0}/2}\frac{1}{\hat{\Gamma}_{B}}z^{-\rho_{X_{T}}}z^{\mu_{X_{T}}}J^{B}$
$=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}H_{B,T,\beta},$
where $\rho_{X_{T}}$ is the operator on $H^{*}_{\mathbb{T}}(B)$ given by multiplication by $c_{1}^{\mathbb{T}}(TB\oplus(E\oplus F^{\vee})^{\oplus r})$ and
$\mu_{X_{T}}(\phi)=\left(\frac{\deg(\phi)}{2}-\frac{\dim(X_{T,\pm})}{2}\right)\phi.$
@@
$\sum_{(e_{1},\ldots,e_{r})\in\mathbb{Z}^{r}}\prod_{k=1}^{r}\prod_{j=1}^{n}\frac{q_{k,+}^{(-y_{k}/2\pi\mathbbm{i}+e_{k})}}{\Gamma(1+(-y_{k}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})\Gamma(1+(y_{k}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}-e_{k})}$
@@
$\sum_{(d_{1},\ldots,d_{r})\in\mathbb{Z}^{r}}\prod_{k=1}^{r}\prod_{j=1}^{n}\frac{q_{k,-}^{(y_{k}/2\pi\mathbbm{i}+d_{k})}}{\Gamma(1+(-y_{k}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d_{k})\Gamma(1+(y_{k}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}+d_{k})}$
@@
(4.2.1) $I_{X_{T,\pm}}(Q,q_{\pm},z)$ $=\psi_{X_{T,\pm}}\left(\sum_{\beta\in NE(B)_{\mathbb{Z}}}\frac{Q^{\beta}}{z^{\beta\cdot c_{1}^{\mathbb{T}}(E\oplus F^{\vee})^{\oplus r}}}H_{B,T,\beta}H_{X_{T,\pm},\beta}\right)$
@@
Let $\mathbb{U}_{H}^{T}:H^{*}_{\mathbb{T}}(X_{T,-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{T,+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ be the map defined by sending $\alpha/e_{\mathbb{T}}(N_{f^{-}})\in H^{*}_{\mathbb{T}}(B_{f^{-}})$ to
$\sum_{f^{+}\in F_{+}}\prod_{i=1}^{r}e^{(n-1)\cdot(x_{f^{+}_{i}}-z_{f^{-}_{i}})/2}\frac{\prod_{j\neq{f^{-}_{i}}}\sin((-z_{j}+x_{f^{+}_{i}})/2\mathbbm{i})}{\prod_{j\neq{f^{-}_{i}}}\sin((-z_{j}+z_{f^{-}_{i}})/2\mathbbm{i})}\frac{\phi_{f^{+},f^{-}}(\alpha)}{e_{\mathbb{T}}(N_{f^{+}})}$
where $\phi_{f^{+},f^{-}}:H^{*}_{\mathbb{T}}(B_{f^{-}})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(B_{f^{+}})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ is the canonical isomorphism induced by the projections $\pi_{f^{+}}:B_{f^{+}}\to B$ and $\pi_{f^{-}}:B_{f^{-}}\to B$, and $N_{f^{\pm}}$ is the normal bundle $N_{B_{f^{\pm}}|X_{\pm}}$.
@@
for all $\beta$, after identifying $q_{k,+}^{-1}=q_{k,-}$. Here the path of analytic continuation is a concatenation
$\gamma_{1}^{T}\star\cdots\star\gamma_{r}^{T}$
where $\gamma_{k}^{T}$ is the path that keeps $q_{i,+}$ constant for $i\neq k$, and in the $\log(q_{k,+})$ plane follows the path $\gamma$ of analytic continuation from the previous section.
Define
$\mathbb{U}^{T}=\psi_{X_{T,+}}\circ\mathbb{U}_{H}^{T}\circ\psi_{X_{T,-}}^{-1}$
in analogy to the previous section. By Theorem 4.2, we obtain:
@@
The linear transformation $\mathbb{U}^{T}$ is symplectic, has a well-defined non-equivariant limit, and satisfies:
@@
5. Grassmann flops
In this section we compute the analytic continuation and symplectic transformation for the $I$-functions $X_{+}$ and $X_{-}$ appearing in the Grassmann flop.
5.1. Analytic continuation
@@
Let $S_{-}$ denote the rank-$r$ tautological bundle on $X_{-}$. We have the exact sequence
@@
and similarly for $X_{+}$. We view $H^{*}_{\mathbb{T}}(X_{\pm})$ as a subspace of $(H^{*}_{\mathbb{T}}(V^{s}(G)/T))$ via Proposition 2.3. After pulling back to $V^{s}(G)/T$, we may write the Gamma class $\hat{\Gamma}_{X_{\pm}}$ as $\hat{\Gamma}_{B}\hat{\Gamma}(TX_{\pm}/B)$ where in this case
$\hat{\Gamma}(TX_{\pm}/B)=\frac{\prod_{k=1}^{r}\prod_{j=1}^{n}\left(\hat{\Gamma}(1+(-y_{k}-x_{j}))\Gamma(1+(y_{k}+z_{j}))\right)}{\hat{\Gamma}(\mathfrak{g}/\mathfrak{t})}$
@@
$I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ $=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(e_{1},\ldots,e_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,+}^{-y_{i}/z+e_{i}}N_{\beta,-e_{1},\ldots,-e_{r}}$
$I_{X_{T,-}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ $=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(d_{1},\ldots,d_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,-}^{y_{i}/z+d_{i}}N_{\beta,d_{1},\ldots,d_{r}}.$
with $N_{\beta,\vec{d}}$ defined in (3.2.5).
@@
$\frac{1}{z^{\beta\cdot c_{1}^{\mathbb{T}}(E\oplus F^{\vee})^{\oplus r}}}\prod_{i\neq k}\frac{\Gamma(1+(y_{k}-y_{i})/z+d_{k}-d_{i})}{\Gamma(1+(y_{k}-y_{i})/z)}\cdot$
$\prod_{k=1}^{r}\prod_{j=1}^{n}\frac{\hat{\Gamma}(1+(-y_{k}-x_{j})/z)\Gamma(1+(y_{k}+z_{j})/z)}{\Gamma(1+(-y_{k}-x_{j})/z+\beta\cdot L_{j}-d_{k})\Gamma(1+(y_{k}+z_{j})/z-\beta\cdot M_{j}+d_{k})},$
@@
$H^{B}$ $:=(2\pi\mathbbm{i})^{-\deg_{0}/2}\frac{1}{\hat{\Gamma}_{B}}z^{-\rho_{X}}z^{\mu_{X}}J^{B}$
$=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}H_{B,\beta},$
where $\rho_{X}$ is the operator on $H^{*}_{\mathbb{T}}(B)$ given by multiplication by $c_{1}^{\mathbb{T}}(TB\oplus(E\oplus F^{\vee})^{\oplus r})$ and $\mu_{X}$ acts on a homogeneous element $\phi$ by
$\mu_{X}(\phi)=\left(\frac{\deg(\phi)}{2}-\frac{\dim(X_{\pm})}{2}\right)\phi.$
@@
$\sum_{(e_{1},\ldots,e_{r})\in\mathbb{Z}^{r}}\prod_{k=1}^{r}\left(q_{k,+}^{(-y_{k}/2\pi\mathbbm{i}+e_{k})}\prod_{i<k}\Gamma(1+(y_{k}-y_{i})/2\pi\mathbbm{i}-e_{k}+e_{i})\Gamma(1-(y_{k}-y_{i})/2\pi\mathbbm{i}+e_{k}-e_{i})\right.$
$\left.\prod_{j=1}^{n}\Gamma(1+(-y_{k}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})^{-1}\Gamma(1+(y_{k}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}-e_{k})^{-1}\right)$
$=$ $\frac{\pi^{r(r-1)}}{\prod_{i<k\leq r}\sin((y_{k}-y_{i})/2\mathbbm{i})}\sum_{(e_{1},\ldots,e_{r})\in\mathbb{Z}^{r}}(-1)^{(r-1)\sum_{k=1}^{r}e_{k}}$
$\prod_{k=1}^{r}\left(q_{k,+}^{(-y_{k}/2\pi\mathbbm{i}+e_{k})}\prod_{i<k}((y_{k}-y_{i})/2\pi\mathbbm{i}-e_{k}+e_{i})\right.$
$\left.\prod_{j=1}^{n}\Gamma(1+(-y_{k}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})^{-1}\Gamma(1+(y_{k}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}-e_{k})^{-1}\right)$
As before we have that
(5.1.1) $I_{X_{+}}(Q,q_{+},z)$ $=\psi_{+}\left(\sum_{\beta\in NE(B)_{\mathbb{Z}}}\frac{Q^{\beta}}{z^{\beta\cdot c_{1}^{\mathbb{T}}(E\oplus F^{\vee})}}H_{B,\beta}H_{X_{+},\beta}\right)\vline_{q_{1,+}=\cdots=q_{r,+}=q_{+}}.$
@@
Restrict $H_{X_{+},\beta}$ to a fixed locus $B_{\delta}$, where $\delta\in D_{+}$ is a size $r$ subset of $\{1,\ldots,n\}$. Then
(5.1.2) $H^{+}_{\beta,\delta}:=H_{X_{+},\beta}|_{B_{\delta}}=\frac{\pi^{r(r-1)}}{\prod_{i<k\leq r}\sin((x_{\delta_{i}}-x_{\delta_{k}})/2\mathbbm{i})}K^{+}_{\beta,\delta}$
@@
$\sum_{e_{1}\geq-\beta\cdot L_{\delta_{1}}}\cdots\sum_{e_{r}\geq-\beta\cdot L_{\delta_{r}}}(-1)^{(r-1)\sum_{k=1}^{r}e_{k}}\prod_{k=1}^{r}\left(q_{k,+}^{(x_{\delta_{k}}/2\pi\mathbbm{i}+e_{k})}\prod_{i<k}((-x_{\delta_{k}}+x_{\delta_{i}})/2\pi\mathbbm{i}-e_{k}+e_{i})\right.$
$\left.\prod_{j=1}^{n}\Gamma(1+(x_{\delta_{k}}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}+e_{k})^{-1}\Gamma(1+(-x_{\delta_{k}}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}-e_{k})^{-1}\right).$
@@
$\sum_{d_{1}\geq\beta\cdot M_{f_{1}}}\cdots\sum_{d_{r}\geq\beta\cdot M_{f_{r}}}(-1)^{(r-1)\sum_{k=1}^{r}d_{k}}\prod_{k=1}^{r}\left(q_{k,-}^{(-z_{f_{k}}/2\pi\mathbbm{i}+d_{k})}\prod_{i<k}((-z_{f_{k}}+z_{f_{i}})/2\pi\mathbbm{i}+d_{k}-d_{i})\right.$
$\left.\prod_{j=1}^{n}\Gamma(1+(z_{f_{k}}-x_{j})/2\pi\mathbbm{i}+\beta\cdot L_{j}-d_{k})^{-1}\Gamma(1+(-z_{f_{k}}+z_{j})/2\pi\mathbbm{i}-\beta\cdot M_{j}+d_{k})^{-1}\right).$
@@
(5.1.3) $H^{-}_{\beta,\delta^{-}}:=H_{X_{-},\beta}|_{B_{\delta^{-}}}=\frac{\pi^{r(r-1)}}{\prod_{i<k\leq r}\sin((z_{\delta^{-}_{i}}-z_{\delta^{-}_{k}})/2\mathbbm{i})}K^{-}_{\beta,\delta^{-}}$
The computation of the analytic continuation of $K_{\beta,\delta}^{+}$ is almost identical to that of $H_{\beta,T,f^{+}}$ from the previous section. In this case the path of analytic continuation will be given by $\gamma_{1}\star\cdots\star\gamma_{r}$ where $\gamma_{k}$ is the path that keeps $q_{i,+}$ constant for $i\neq k$ and in the $\log(q_{k,+})$ plane, will go through $\mathfrak{Im}(\log(q_{k,+}))=(n-r)\pi\mathbbm{i}$ when $\mathfrak{R}(\log(q_{k,+}))$ is close to 0. See Appendix A for details.
@@
$C^{K}_{f^{-},f^{+}}=\prod_{i=1}^{r}e^{(n-r)\cdot(x_{f^{+}_{i}}-z_{f^{-}_{i}})/2}\frac{\prod_{j\neq{f^{-}_{i}}}\sin((-z_{j}+x_{f^{+}_{i}})/2\mathbbm{i})}{\prod_{j\neq{f^{-}_{i}}}\sin((-z_{j}+z_{f^{-}_{i}})/2\mathbbm{i})}$
@@
(5.1.4) $\sum_{f^{-}\in F_{-}}C^{K}_{f^{-},\delta^{+}}K^{-}_{\beta,f^{-}},$
after identifying $q_{k,+}^{-1}=q_{k,-}$. We emphasize that here the sum is over all functions $f^{-}:\{1,\ldots,r\}\to\{1,\ldots,n\}$.
Then by (5.1.2) and (5.1.3), if we define
$C^{H}_{f^{-},\delta^{+}}=C^{K}_{f^{-},\delta^{+}}\prod_{i<k\leq r}\frac{\sin((z_{f^{-}_{i}}-z_{f^{-}_{k}})/2\mathbbm{i})}{\sin((x_{\delta^{+}_{i}}-x_{\delta^{+}_{k}})/2\mathbbm{i})},$
@@
Note that the numerator of $C^{H}_{f^{-},\delta^{+}}$ vanishes unless $f^{-}_{i}\neq f^{-}_{k}$ for $i\neq k$. Therefore we can replace the summation above to be only over all injective functions $\delta^{-}\in In_{-}$:
(5.1.5) $\widetilde{H^{+}_{\beta,\delta^{+}}}=\sum_{\delta^{-}\in In_{-}}C^{H}_{\delta^{-},\delta^{+}}H^{-}_{\beta,\delta^{-}}.$
In order to connect the above computation to the actual $I$-functions for $X_{+}$ and $X_{-}$ via (5.1.1), we must homotope the path $\gamma_{1}\star\cdots\star\gamma_{r}$ of analytic continuation to a new path which lies entirely within the diagonal $q_{1,+}=\cdots=q_{r,+}$.
@@
The path of analytic continuation $\gamma_{1}\star\cdots\star\gamma_{r}$ used in (5.1.5) is homotopic to a path $\hat{\gamma}$ contained entirely in the locus $q_{1,+}=\cdots=q_{r,+}$, via a homotopy which is
  1. (1)

### v2

@@ 4.2 Products of projective bundles
In this case, for any $j$ we have the relationship between Gamma functions:
@@
$\hat{\Gamma}(TX_{T}/B)=\prod_{k=1}^{r}\prod_{j=1}^{n}\bigl{(}\hat{\Gamma}(1+(-y_{k}-x_{j}))\Gamma(1+(y_{k}+z_{j}))\bigr{)}.$
@@
$H^{B}_{T}:=(2\pi{\rm i})^{-\deg_{0}/2}\frac{1}{\hat{\Gamma}_{B}}z^{-\rho_{X_{T}}}z^{\mu_{X_{T}}}J^{B}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}H_{B,T,\beta},$
where $\rho_{X_{T}}$ is the operator on $H^{*}_{\mathbb{T}}(B)$ given by multiplication by $c_{1}^{\mathbb{T}}\bigl{(}TB\oplus\bigl{(}E\oplus F^{\vee}\bigr{)}^{\oplus r}\bigr{)}$ and
$\mu_{X_{T}}(\phi)=\biggl{(}\frac{\deg(\phi)}{2}-\frac{\dim(X_{T,\pm})}{2}\biggr{)}\phi.$
@@
$\sum_{(e_{1},\dots,e_{r})\in\mathbb{Z}^{r}}\prod_{k=1}^{r}\prod_{j=1}^{n}\!\frac{q_{k,+}^{(-y_{k}/2\pi{\rm i}+e_{k})}}{\Gamma(1+(-y_{k}-x_{j})/2\pi{\rm i}+\beta\cdot L_{j}+e_{k})\Gamma(1+(y_{k}+z_{j})/2\pi{\rm i}-\beta\cdot M_{j}-e_{k})},$
@@
$\sum_{(d_{1},\dots,d_{r})\in\mathbb{Z}^{r}}\prod_{k=1}^{r}\prod_{j=1}^{n}\!\frac{q_{k,-}^{(y_{k}/2\pi{\rm i}+d_{k})}}{\Gamma(1+(-y_{k}-x_{j})/2\pi{\rm i}+\beta\cdot L_{j}-d_{k})\Gamma(1+(y_{k}+z_{j})/2\pi{\rm i}-\beta\cdot M_{j}+d_{k})}.$
@@
$I_{X_{T,\pm}}(Q,q_{\pm},z)$ ${}=\psi_{X_{T,\pm}}\biggl{(}\sum_{\beta\in NE(B)_{\mathbb{Z}}}\frac{Q^{\beta}}{z^{\beta\cdot c_{1}^{\mathbb{T}}(E\oplus F^{\vee})^{\oplus r}}}H_{B,T,\beta}H_{X_{T,\pm},\beta}\biggr{)}.$
@@
Let $\mathbb{U}_{H}^{T}\colon H^{*}_{\mathbb{T}}(X_{T,-})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}(X_{T,+})\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ be the map defined by sending $\alpha/e_{\mathbb{T}}\bigl{(}N_{f^{-}}\bigr{)}\in H^{*}_{\mathbb{T}}\bigl{(}B_{f^{-}}\bigr{)}$ to
$\sum_{f^{+}\in F_{+}}\prod_{i=1}^{r}{\rm e}^{(n-1)\cdot(x_{f^{+}_{i}}-z_{f^{-}_{i}})/2}\frac{\prod_{j\neq{f^{-}_{i}}}\sin\bigl{(}\bigl{(}-z_{j}+x_{f^{+}_{i}}\bigr{)}/2{\rm i}\bigr{)}}{\prod_{j\neq{f^{-}_{i}}}\sin\bigl{(}\bigl{(}-z_{j}+z_{f^{-}_{i}}\bigr{)}/2{\rm i}\bigr{)}}\frac{\phi_{f^{+},f^{-}}(\alpha)}{e_{\mathbb{T}}(N_{f^{+}})},$
where $\phi_{f^{+},f^{-}}\colon H^{*}_{\mathbb{T}}\bigl{(}B_{f^{-}}\bigr{)}\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}\to H^{*}_{\mathbb{T}}\bigl{(}B_{f^{+}}\bigr{)}\otimes_{R_{\mathbb{T}}}S_{\mathbb{T}}$ is the canonical isomorphism induced by the projections $\pi_{f^{+}}\colon B_{f^{+}}\to B$ and $\pi_{f^{-}}\colon B_{f^{-}}\to B$, and $N_{f^{\pm}}$ is the normal bundle $N_{B_{f^{\pm}}|X_{\pm}}$.
@@
for all $\beta$, after identifying $q_{k,+}^{-1}=q_{k,-}$. Here the path of analytic continuation is a concatenation $\gamma_{1}^{T}\star\cdots\star\gamma_{r}^{T}$, where $\gamma_{k}^{T}$ is the path that keeps $q_{i,+}$ constant for $i\neq k$, and in the $\log(q_{k,+})$ plane follows the path $\gamma$ of analytic continuation from the previous section.
Define $\mathbb{U}^{T}=\psi_{X_{T,+}}\circ\mathbb{U}_{H}^{T}\circ\psi_{X_{T,-}}^{-1}$ in analogy to the previous section. By Theorem 4.2, we obtain the following.
@@
The linear transformation $\mathbb{U}^{T}$ is symplectic, has a well-defined non-equivariant limit, and satisfies
@@
5 Grassmann flops
In this section, we compute the analytic continuation and symplectic transformation for the $I$-functions $X_{+}$ and $X_{-}$ appearing in the Grassmann flop.
5.1 Analytic continuation
@@
Let $S_{-}$ denote the rank-$r$ tautological bundle on $X_{-}$. By (2.5), we have the short exact sequence
@@
and similarly for $X_{+}$. We view $H^{*}_{\mathbb{T}}(X_{\pm})$ as a subspace of $\bigl{(}H^{*}_{\mathbb{T}}(V^{s}(G)/T)\bigr{)}$ via Proposition 2.3. After pulling back to $V^{s}(G)/T$, we may write the Gamma class $\hat{\Gamma}_{X_{\pm}}$ as $\hat{\Gamma}_{B}\hat{\Gamma}(TX_{\pm}/B)$, where in this case
$\hat{\Gamma}(TX_{\pm}/B)=\frac{\prod_{k=1}^{r}\prod_{j=1}^{n}\bigl{(}\hat{\Gamma}(1+(-y_{k}-x_{j}))\Gamma(1+(y_{k}+z_{j}))\bigr{)}}{\hat{\Gamma}(\mathfrak{g}/\mathfrak{t})}$
@@
$I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(e_{1},\dots,e_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,+}^{-y_{i}/z+e_{i}}N_{\beta,-e_{1},\dots,-e_{r}},$
$I_{X_{T,-}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{(d_{1},\dots,d_{r})\in\mathbb{Z}^{r}}\prod_{i=1}^{r}q_{i,-}^{y_{i}/z+d_{i}}N_{\beta,d_{1},\dots,d_{r}}$
with $N_{\beta,\vec{d}}$ defined in (3.4).
@@
$\frac{1}{z^{\beta\cdot c_{1}^{\mathbb{T}}(E\oplus F^{\vee})^{\oplus r}}}\prod_{i\neq k}\frac{\Gamma(1+(y_{k}-y_{i})/z+d_{k}-d_{i})}{\Gamma(1+(y_{k}-y_{i})/z)}$
×$\qquad{}\times\prod_{k=1}^{r}\prod_{j=1}^{n}\frac{\hat{\Gamma}(1+(-y_{k}-x_{j})/z)\Gamma(1+(y_{k}+z_{j})/z)}{\Gamma(1+(-y_{k}-x_{j})/z+\beta\cdot L_{j}-d_{k})\Gamma(1+(y_{k}+z_{j})/z-\beta\cdot M_{j}+d_{k})}.$
@@
$H^{B}:=(2\pi{\rm i})^{-\deg_{0}/2}\frac{1}{\hat{\Gamma}_{B}}z^{-\rho_{X}}z^{\mu_{X}}J^{B}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}H_{B,\beta},$
where $\rho_{X}$ is the operator on $H^{*}_{\mathbb{T}}(B)$ given by multiplication by $c_{1}^{\mathbb{T}}\bigl{(}TB\oplus\bigl{(}E\oplus F^{\vee}\bigr{)}^{\oplus r}\bigr{)}$ and $\mu_{X}$ acts on a homogeneous element $\phi$ by
$\mu_{X}(\phi)=\biggl{(}\frac{\deg(\phi)}{2}-\frac{\dim(X_{\pm})}{2}\biggr{)}\phi.$
@@
$\sum_{(e_{1},\dots,e_{r})\in\mathbb{Z}^{r}}\prod_{k=1}^{r}\!\Biggl{(}\!q_{k,+}^{(-y_{k}/2\pi{\rm i}+e_{k})}\!\prod_{i<k}\!\Gamma(1+(y_{k}-y_{i})/2\pi{\rm i}-e_{k}+e_{i})\Gamma(1-(y_{k}-y_{i})/2\pi{\rm i}+e_{k}-e_{i})\!$
×$\quad\quad{}\times\prod_{j=1}^{n}\Gamma(1+(-y_{k}-x_{j})/2\pi{\rm i}+\beta\cdot L_{j}+e_{k})^{-1}\Gamma(1+(y_{k}+z_{j})/2\pi{\rm i}-\beta\cdot M_{j}-e_{k})^{-1}\!\Biggr{)}$
$\quad{}=\frac{\pi^{r(r-1)/2}}{\prod_{i<k\leq r}\sin((y_{k}-y_{i})/2{\rm i})}\sum_{(e_{1},\dots,e_{r})\in\mathbb{Z}^{r}}(-1)^{(r-1)\sum_{k=1}^{r}e_{k}}$
×$\qquad{}\times\prod_{k=1}^{r}\Biggl{(}q_{k,+}^{(-y_{k}/2\pi{\rm i}+e_{k})}\prod_{i<k}((y_{k}-y_{i})/2\pi{\rm i}-e_{k}+e_{i})$
×$\qquad{}\times\prod_{j=1}^{n}\Gamma(1+(-y_{k}-x_{j})/2\pi{\rm i}+\beta\cdot L_{j}+e_{k})^{-1}\Gamma(1+(y_{k}+z_{j})/2\pi{\rm i}-\beta\cdot M_{j}-e_{k})^{-1}\Biggr{)}.$
As before, we have that
$I_{X_{+}}(Q,q_{+},z)$ ${}=\psi_{+}\left(\sum_{\beta\in NE(B)_{\mathbb{Z}}}\frac{Q^{\beta}}{z^{\beta\cdot c_{1}^{\mathbb{T}}(E\oplus F^{\vee})}}H_{B,\beta}H_{X_{+},\beta}\right)\vline_{q_{1,+}=\cdots=q_{r,+}=q_{+}}.$ (5.1)
@@
Restrict $H_{X_{+},\beta}$ to a fixed locus $B_{\delta}$, where $\delta\in D_{+}$ is a size $r$ subset of $\{1,\dots,n\}$. Then
$H^{+}_{\beta,\delta}:=H_{X_{+},\beta}|_{B_{\delta}}=\frac{\pi^{r(r-1)/2}}{\prod_{i<k\leq r}\sin((x_{\delta_{i}}-x_{\delta_{k}})/2{\rm i})}K^{+}_{\beta,\delta},$ (5.2)
@@
$\sum_{e_{1}\geq-\beta\cdot L_{\delta_{1}}}\cdots\sum_{e_{r}\geq-\beta\cdot L_{\delta_{r}}}(-1)^{(r-1)\sum_{k=1}^{r}e_{k}}\prod_{k=1}^{r}\Biggl{(}q_{k,+}^{(x_{\delta_{k}}/2\pi{\rm i}+e_{k})}\prod_{i<k}((-x_{\delta_{k}}+x_{\delta_{i}})/2\pi{\rm i}-e_{k}+e_{i})$
×$\quad{}\times\prod_{j=1}^{n}\Gamma(1+(x_{\delta_{k}}-x_{j})/2\pi{\rm i}+\beta\cdot L_{j}+e_{k})^{-1}\Gamma(1+(-x_{\delta_{k}}+z_{j})/2\pi{\rm i}-\beta\cdot M_{j}-e_{k})^{-1}\Biggr{)}.$
@@
$\sum_{d_{1}\geq\beta\cdot M_{f_{1}}}\cdots\sum_{d_{r}\geq\beta\cdot M_{f_{r}}}(-1)^{(r-1)\sum_{k=1}^{r}d_{k}}\prod_{k=1}^{r}\Biggl{(}q_{k,-}^{(-z_{f_{k}}/2\pi{\rm i}+d_{k})}\prod_{i<k}((-z_{f_{k}}+z_{f_{i}})/2\pi{\rm i}+d_{k}-d_{i})$
×$\quad{}\times\prod_{j=1}^{n}\Gamma(1+(z_{f_{k}}-x_{j})/2\pi{\rm i}+\beta\cdot L_{j}-d_{k})^{-1}\Gamma(1+(-z_{f_{k}}+z_{j})/2\pi{\rm i}-\beta\cdot M_{j}+d_{k})^{-1}\Biggr{)}.$
@@
$H^{-}_{\beta,\delta^{-}}:=H_{X_{-},\beta}|_{B_{\delta^{-}}}=\frac{\pi^{r(r-1)/2}}{\prod_{i<k\leq r}\sin\bigl{(}\bigl{(}z_{\delta^{-}_{i}}-z_{\delta^{-}_{k}}\bigr{)}/2{\rm i}\bigr{)}}K^{-}_{\beta,\delta^{-}}.$ (5.3)
The computation of the analytic continuation of $K_{\beta,\delta}^{+}$ is almost identical to that of $H_{\beta,T,f^{+}}$ from the previous section. In this case the path of analytic continuation will be given by $\gamma_{1}\star\cdots\star\gamma_{r}$ where $\gamma_{k}$ is the path that keeps $q_{i,+}$ constant for $i\neq k$ and in the $\log(q_{k,+})$ plane, will go through $\mathfrak{Im}(\log(q_{k,+}))=(n-r)\pi{\rm i}$ when $\mathfrak{Re}(\log(q_{k,+}))$ is close to 0. See Appendix A for details.
@@
$C^{K}_{f^{-},f^{+}}=\prod_{i=1}^{r}{\rm e}^{(n-r)\cdot(x_{f^{+}_{i}}-z_{f^{-}_{i}})/2}\frac{\prod_{j\neq{f^{-}_{i}}}\sin\bigl{(}\bigl{(}-z_{j}+x_{f^{+}_{i}}\bigr{)}/2{\rm i}\bigr{)}}{\prod_{j\neq{f^{-}_{i}}}\sin\bigl{(}\bigl{(}-z_{j}+z_{f^{-}_{i}}\bigr{)}/2{\rm i}\bigr{)}}.$
@@
$\sum_{f^{-}\in F_{-}}C^{K}_{f^{-},\delta^{+}}K^{-}_{\beta,f^{-}},$ (5.4)
after identifying $q_{k,+}^{-1}=q_{k,-}$. We emphasize that here the sum is over all functions $f^{-}\colon\{1,\dots,r\}\allowbreak\to\{1,\dots,n\}$.
Then by (5.2) and (5.3), if we define
$C^{H}_{f^{-},\delta^{+}}=C^{K}_{f^{-},\delta^{+}}\prod_{i<k\leq r}\frac{\sin\bigl{(}\bigl{(}z_{f^{-}_{i}}-z_{f^{-}_{k}}\bigr{)}/2{\rm i}\bigr{)}}{\sin\bigl{(}\bigl{(}x_{\delta^{+}_{i}}-x_{\delta^{+}_{k}}\bigr{)}/2{\rm i}\bigr{)}},$
@@
Note that the numerator of $C^{H}_{f^{-},\delta^{+}}$ vanishes unless $f^{-}_{i}\neq f^{-}_{k}$ for $i\neq k$. Therefore, we can replace the summation above to be only over all injective functions $\delta^{-}\in{\rm In}_{-}$:
$\widetilde{H^{+}_{\beta,\delta^{+}}}=\sum_{\delta^{-}\in{\rm In}_{-}}C^{H}_{\delta^{-},\delta^{+}}H^{-}_{\beta,\delta^{-}}.$ (5.5)
In order to connect the above computation to the actual $I$-functions for $X_{+}$ and $X_{-}$ via (5.1), we must homotope the path $\gamma_{1}\star\cdots\star\gamma_{r}$ of analytic continuation to a new path which lies entirely within the diagonal $q_{1,+}=\cdots=q_{r,+}$.
@@
The path of analytic continuation $\gamma_{1}\star\cdots\star\gamma_{r}$ used in (5.5) is homotopic to a path $\hat{\gamma}$ contained entirely in the locus $q_{1,+}=\cdots=q_{r,+}$, via a homotopy which is
  1. 1)1)1 )